OpenStudy (anonymous):

If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and second equations, then the second and third equations. 3x – 5y + 2z = 7 4x + 9y – z = 2 x – 3y + 3z = 7 If you were to use the elimination method to solve the following system, choose the new system of equations that would result after the variable z is eliminated in the first and second equations, then the second and third equations. 3x – 5y + 2z = 7 4x + 9y – z = 2 x – 3y + 3z = 7 @Mathematics

OpenStudy (anonymous):

11x + 13y = 11 3x – 5y = 7 x – 3y = 7 13x + 24y = 13 x – 3y = 7 3x – 5y = 7 11x + 13y = 11 13x + 24y = 13